Determine the area of the given curves


1 . 𝑦2 = 2 π‘₯ ; 𝑦 = π‘₯ βˆ’ 4


2. 𝑦 = π‘₯ βˆ’ π‘₯2 ; 𝑦 = βˆ’ π‘₯


1
Expert's answer
2022-05-20T07:16:37-0400

1)  𝑦2Β = 2 π‘₯ ; 𝑦 = π‘₯ βˆ’ 4

S="\int"-24(y+4-y2/2)dy=(y2/2+4y-y3/6)"\mid"-24=8+16-64/6-2+8-8/6=30-32/3-4/3=30-36/3=30-12=18



2) 𝑦 = π‘₯ βˆ’ π‘₯2Β ; 𝑦 = βˆ’ π‘₯

S="\int"02(x-x2+x)dx="\int"02(2x-x2) dx=(x2-x3/3)"\mid"02=4-8/3=4/3